OpenStudy (anonymous):

Which is a solution to the equation y=5x-6 A. (3, 21) B. (4, 2) C. (5, 35) D. (6, 24)

OpenStudy (anonymous):

substitute the values in to the equation. for example: (3,21) 21 = 5(3)-6 21 = 15 - 6 21 = 9 no, so A is not a solution. Do this with the remaining choices